# Faraday Depolarization Study of a Radio Galaxy Using LOFAR Two-metre Sky Survey: Data Release 2

### v1

@@ Abstract
We present a detailed depolarization analysis of a radio galaxy ILTJ012215.21+254334.8, utilizing polarimetric data from the LOFAR Two-metre Sky Survey (LoTSS) Data Release 2 (DR2) catalogue. The selected source exhibits a rotation measure (RM) of rad m-2 and a projected linear size of kpc at a redshift . Depolarization model fitting was performed on LOFAR High Band Antenna data (120–168 MHz), with fractional polarization detected at . Five depolarization models were tested, and Bayesian qu-fitting revealed that the three-component model (1T+2ED) best describes the data, with a reduced chi-squared value of 2.12 and a logarithmic Bayesian evidence of 1384.82. This model includes a Faraday-thin component at RM rad m-2 (instrumental leakage) and two external Faraday dispersion astrophysical emission at RM rad m-2. The results demonstrate that depolarization in low-frequency radio galaxies requires multi-component modelling and is driven by turbulence and inhomogeneity in the magneto-ionic medium. Our findings highlight the potential of LOFAR polarization studies for probing galactic and intergalactic magnetic fields with high precision.
@@
techniques: polarimetric, galaxies: active, galaxies: magnetic fields, radio continuum: galaxies, radiation mechanisms: non-thermal1 Introduction
@@
Radio galaxies exhibit a wide range of physical sizes, from a few parsecs up to several megaparsecs, and are morphologically classified according to the Fanaroff-Riley (FR) scheme (Fanaroff and Riley, 1974). FR II galaxies are powerful, edge-brightened sources with luminous hotspots at the extremities of their lobes. They commonly show weak but well-collimated jets that often appear one-sided owing to relativistic beaming, with radio luminosities typically above at 178 MHz (Fanaroff and Riley, 1974). These systems are typically associated with quasars or more distant radio-loud galaxies. Conversely, FR I galaxies are weaker, edge-darkened sources whose emission is brightest near the core and decreases outward. They generally lack prominent hotspots, and their jets and lobes often appear distorted. With radio luminosities below , FR I sources are more frequently found in nearby galaxies or dense galaxy clusters environment. Although, Mingo et al. (2019) recently confirmed for LoTSS catalogue radio luminosity is not always a reliable way to determine if a source is edge-brightened (FR-II) or centre-brightened (FR-I).
@@
In this paper, we investigate the depolarization properties of a single source drawn from the LoTSS DR2 catalogue (O’Sullivan et al., 2023). Located at a redshift of 0.05, with a radio luminosity of at 144 MHz, this source can be a FR I radio galaxy (Fanaroff and Riley, 1974). We employ the qu-fitting technique to model the fractional Stokes parameters as functions of wavelength squared. Recent studies have demonstrated that this approach is often more sensitive than RM synthesis for detecting multiple Faraday components (Farnsworth et al., 2011; O’Sullivan et al., 2012; Sun et al., 2015).
The study of Faraday depolarization in radio galaxies provides a critical window into the magneto-ionic environments both within and surrounding these powerful extragalactic sources (Strom, 1973). Leveraging the LOFAR Two-metre Sky Survey Data Release 2 (LoTSS-DR2), which offers high-sensitivity, high-resolution polarization data at low radio frequencies enables unprecedented insight into the interplay between magnetic fields, plasma conditions, and large-scale radio structures. The paper is organized as follows: Section 2 introduces the key observables of radio polarization and Faraday rotation; Section 3 describes the depolarization models applied in our analysis; Section 4 outlines the main properties of the selected LoTSS DR2 source; Section 5 details the methodology; Section 6 presents the results; and Section 7 discusses the implications and provides prospects for future work.
@@ 2 Polarization and Faraday Rotation
Polarization in radio galaxies offers direct evidence for synchrotron radiation, caused by relativistic electrons spiralling along magnetic fields—which naturally produces linearly polarized light. Measuring the degree and orientation of polarization reveals the strength and structure of magnetic fields in jets and lobes, as well as information about the galaxy’s orientation relative to the observer.To quantitatively describe the polarization state of electromagnetic radiation, a set of four values is utilized known as Stokes parameters: I, Q, U, and V. These parameters are experimentally convenient as they correspond to sums or differences of measurable intensities, allowing for direct determination from observations. The complex linear polarization is defined as:
@@
The study of the polarization states of radio emissions enables us to analyse two important effects: Faraday Rotation and Faraday Depolarization (Anderson et al., 2015; Piras, 2024). Faraday Rotation is caused by the rotation of the plane of polarization of electromagnetic waves as they journey across vast cosmic distances interacting with magnetized medium. Crucially, the extent of this rotation directly encodes vital information about the magneto-ionic medium—the ionized gas and embedded magnetic fields. This makes Faraday rotation an exceptionally powerful tool for probing magnetic fields not just within galaxies, but also in the vast and tenuous intergalactic medium, providing unique insights into the large-scale magnetization of the Universe that cannot be obtained through other means (Brentjens and de Bruyn, 2005). To accurately characterize the Faraday rotation for these distinct regions of polarized emission, the concept of Faraday Depth is utilized. It is specifically defined as the integral of the thermal electron density (in units of )and the line-of-sight component of the magnetic field (in ) over the physical distance (x) (in parsec) to the source, typically expressed as (Pasetto et al., 2018) :
In the most straightforward scenario, where a polarized radio source lies behind a single, uniform layer of magnetized plasma, the concept of Faraday depth simplifies considerably: it becomes equivalent to the Rotation Measure (RM) itself. In such an ideal case, the observed polarization angle is directly determined by the source’s intrinsic polarization angle and the Faraday rotation (RM), following the simple linear relationship (Anderson et al., 2016):
@@ 3 Depolarization
Depolarization occurs because variations in the magnetic field or electron density along the line of sight, or across the telescope beam, cause the polarization angle to rotate by different amounts in different regions, reducing the net observed polarization (Burn, 1966; Goodlet et al., 2004). When these diverse polarization orientations are averaged across the observing beam, they tend to cancel each other out, diminishing the overall polarization signal. Hence, two types of depolarization come into play: internal and external (Pasetto et al., 2018). The remaining polarization percentage is effectively utilized to gauge out the degree of order and orientation within the magnetic field, whether it resides in the radio source itself or in the intervening inter galactic medium (IGM). The complex polarized signal observed from radio sources is fundamentally described by a model that accounts only for Faraday rotation and does not depolarize the signal is given as (Burn, 1966) :
Here, represents the intrinsic fractional polarization of the radiation, and is the intrinsic polarization angle at the source of the emission. The term describes the Faraday rotation caused by the foreground magneto-ionic medium, where RM is the Rotation Measure and is the square of the observing wavelength. This model is also known as ‘Faraday Thin’ Component (Burn, 1966). In real scenarios, the magnetic field is not perfectly ordered, there exists random and uniform fields causing depolarization (Goodlet et al., 2004). The various depolarization contributions are thus added to this fundamental equation 5.
@@ 4 The LoTSS DR2 Catalogue and the Selected Source
In this pilot study, we focus our depolarization modelling on a source, ILTJ012215.21+25433.8, with an estimated luminosity of W Hz-1 at 144 MHz, RM of rad m-2, and an estimated projected linear size of 334.94 kpc (O’Sullivan et al., 2023). A relatively larger projected linear size in radio galaxies like ILTJ012215.21+254334.8 enables more reliable and physically informative depolarization modelling, as the increased path length through the magnetized plasma allows more lines of sight for Faraday rotation and depolarization effects to manifest. Furthermore, the source’s projected linear size is close to the LoTSS-DR2 median linear size of kpc (O’Sullivan et al., 2023), providing an opportunity to investigate depolarization properties of a typical galaxy in the LoTSS-DR2 catalogue. In a subsequent study, we plan to investigate the depolarization properties of all FR-I samples in the DR2 catalogue.
@@
Here, RM is the Rotation Measure; P is the linear polarization; is the fractional polarization; I is the total intensity; is the luminosity at 144 MHz; is the projected linear size; and z is the redshift.
@@ 5.2 The priors
In qu-fitting, setting good starting assumptions (priors) for the model parameters is important because they help guide the fitting process in the right direction. The selection of these priors is often informed by the observed Faraday spectrum of the source (Piras, 2024). For example, when combining multiple models, the sum of the initial polarisation percentages for all components must not exceed 1, and each initial polarization angle is limited between 0 and 180 degrees for all the models considered in Table 2.
@@
To analyse the observed Stokes and values, along with their errors ( and ), a one-dimensional qu-fitting technique was applied to the LoTSS DR2 polarization spectra of the source111https://lofar-mksp.org/data/. The Stokes spectral behaviour for a 97.6 kHz channel bandwidth is unreliable and thus not included in the spectra (O’Sullivan et al., 2023). However, we generate model Stokes values using the 144 MHz value, assuming a spectral index of -0.7 for the source. For the Stokes error (), we estimate the uncertainty across frequency by assuming random variations on the single-frequency catalogue value. These values are then used to create the fractional and spectra. The corresponding and errors ( and ) are calculated as,
We used the RM-TOOLS software222https://github.com/CIRADA-Tools/RM-Tools/tree/master (Purcell et al., 2020), which incorporates the MultiNest algorithm (Feroz and Hobson, 2008; Feroz et al., 2009, 2019). MultiNest, implemented via PYMULTINEST (Buchner, J. et al., 2014), is employed to efficiently explore the parameter space of the chosen model, identify the best-fit parameters, and estimate their associated uncertainties. The algorithm provides the posterior distribution for the model parameters and the Bayesian evidence () for the model, which quantifies the probability of observing the data given a specific model.
@@ 6 Results
The comparison of different models based on the odds ratio and the reduced chi-squared statistic () indicate that the best description of the depolarization behaviour of the source is provided by the 1T+2ED model. This particular model includes three polarization components: one Faraday-thin term corresponding to instrumental leakage, located at a rotation measure (RM) of about , and two external depolarization (2ED) components, centred near , which likely represent the astrophysical polarized emission (O’Sullivan et al., 2023). The astrophysical emission is better represented by the superposition of two external depolarization components at similar RMs, each capturing slightly different aspects of turbulence or inhomogeneities in the magneto-ionic medium along the line of sight (Sokoloff et al., 1998; O’Sullivan et al., 2017).
@@
| 1T+2ED | 11 | 3 | 814 | 2.12 | 1384.82 | 0 |
| 1Tk+2ED | 11 | 3 | 814 | 2.12 | 1376.30 | 17.05 |
| 1T+1ED | 7 | 2 | 818 | 2.77 | 1154.81 | 460.04 |
| 1Tk+1ED | 7 | 2 | 818 | 2.77 | 1148.50 | 472.65 |
| 1ED | 4 | 1 | 821 | 2.94 | 1097.36 | 574.93 |
@@
Figure 2 shows the polarization angle as a function of wavelength squared. This plot reveals the rotation of the plane of polarization due to the Faraday effect. We observe an overall linear increase or decrease of with , indicating a coherent Faraday rotation caused by a relatively uniform magnetized plasma along the line of sight. However, the best-fit model also indicates some flattening or irregular fluctuations around , suggesting that partial depolarization is occurring around these wavelengths.
Figure 3 presents the polarization vector in the complex plane, where the horizontal axis is q and the vertical axis is u. Ideally, without depolarization, the data would trace a circular arc of constant radius as changes. However, in the observed data the radius shrinks, showing that the polarization fraction decreases with increasing . This contraction of the polarization vector’s length is direct evidence of depolarization, and the best-fit model captures this trend adequately. Moreover, higher frequency data (red) near the centre suggest strong and coherent polarization, while lighter points (violet/blue) that spread outward indicate depolarization and more complex polarization behaviour at lower frequencies. This pattern is characteristic of Faraday depolarization in polarized radio sources. Finally, the posterior distribution of the best-fit model parameters along with error bars is shown in Figure 4. Note that the derived values of in Table 4 are not meaningful due to the lack of absolute calibration for the polarization angles.

### v2

@@ Abstract
We present a detailed depolarization analysis of the radio galaxy ILTJ012215.21+254334.8 using polarimetric data from the LOFAR Two-metre Sky Survey (LoTSS) Data Release 2 (DR2) catalogue. This source, with RM 47 rad m-2 and projected linear size 335 kpc at , serves as a pilot for systematic QU-fitting of unresolved LoTSS sources, building on prior single-target studies that established the need for multi-component Faraday models in complex magneto-ionic media.
Fitting five depolarization models to the LoTSS HBA (120-168 MHz) fractional polarization spectra reveals a decisively preferred three-component model (one Faraday-thin instrumental leakage, plus two external Faraday dispersions), demonstrating that LoTSS data alone can constrain moderate Faraday complexity in typical DR2 galaxies. Our results highlight turbulence and inhomogeneity in the foreground magneto-ionic medium and outline a path for population studies of LoTSS FR-I sources.
@@
techniques: polarimetric, galaxies: active, galaxies: magnetic fields, radio : galaxies, radiation mechanisms: non-thermal1 Introduction
@@
Radio galaxies exhibit a wide range of physical sizes, from a few parsecs up to several megaparsecs, and are morphologically classified according to the Fanaroff-Riley (FR) scheme (Fanaroff and Riley, 1974). FR II galaxies are powerful, edge-brightened sources with luminous hotspots at the extremities of their lobes. They commonly show weak but well-collimated jets that often appear one-sided owing to relativistic beaming, with radio luminosities typically above at 178 MHz (Fanaroff and Riley, 1974). These systems are typically associated with quasars or more distant radio-loud galaxies. Conversely, FR I galaxies are weaker, edge-darkened sources whose emission is brightest near the core and decreases outward. They generally lack prominent hotspots, and their jets and lobes often appear distorted. With radio luminosities below , FR I sources are more frequently found in nearby galaxies or dense galaxy clusters environment. Although Mingo et al. (2019) recently confirmed for LoTSS catalogue radio luminosity is not always a reliable way to determine if a source is edge-brightened (FR-II) or centre-brightened (FR-I).
@@
Broadband QU-fitting of individual radio galaxies has already established that multiple Faraday components are often required to describe their polarization behaviour, and that magneto-ionic environments are typically inhomogeneous and turbulent (Farnsworth et al., 2011; Anderson et al., 2015; O’Sullivan et al., 2017). Such work has largely focused either on bright, well-studied individual AGN with wide bandwidth coverage, or on small samples selected for particular environments (e.g., cluster cores, remnant sources). At higher angular resolution, resolved polarimetric imaging has revealed strong spatial variations in RM and depolarization across jets and lobes, further highlighting the complexity of the underlying magnetic field and thermal gas distributions (O’Sullivan et al., 2012; Pasetto et al., 2018). However, there has been relatively little effort to connect such detailed studies to the emerging large low-frequency RM grids such as LoTSS-DR2, which now provide homogeneous polarization measurements for thousands of sources (Oei et al., 2023).
The specific aim here is to use LOFAR HBA data for one representative LoTSS-DR2 source to test how well standard depolarization models can be constrained from unresolved fractional Stokes spectra alone, and to characterize the multi-component Faraday structure implied by the LoTSS-only data. This single-source analysis is designed as a proof of concept for a future systematic qu-fitting study of the FR-I subsample in LoTSS-DR2, where the same methodology can be applied uniformly to a statistically meaningful population. Here, we investigate the depolarization properties of a single source drawn from the LoTSS DR2 catalogue (O’Sullivan et al., 2023). Located at a redshift of 0.05, with a radio luminosity of at 144 MHz, this source can be a FR I radio galaxy (Fanaroff and Riley, 1974). We employ the qu-fitting technique to model the fractional Stokes parameters as functions of wavelength squared. Recent studies have demonstrated that this approach is often more sensitive than RM synthesis for detecting multiple Faraday components (Farnsworth et al., 2011; O’Sullivan et al., 2012; Sun et al., 2015).
The study of Faraday depolarization in radio galaxies provides a critical window into the magneto-ionic environments both within and surrounding these powerful extragalactic sources (Strom, 1973). The goal of this study is to analyze LOFAR HBA data for a representative source from the LoTSS-DR2 catalog to test how well standard depolarization models can be constrained from unresolved fractional Stokes spectra alone, and to characterize the multi-component Faraday structure implied by the LoTSS-only data. This single-source analysis is designed as a proof of concept for a future systematic qu-fitting study of the FR-I subsample in LoTSS-DR2, where the same methodology can be applied uniformly to a statistically meaningful population. The paper is organized as follows: Section 2 introduces the key observables of radio polarization and Faraday rotation; Section 3 describes the depolarization models applied in our analysis; Section 4 outlines the main properties of the selected LoTSS DR2 source; Section 5 details the methodology; Section 6 presents the results; and Section 7 discusses the implications and provides prospects for future work.
@@ 2 Polarization and Faraday Rotation
Polarization in radio galaxies offers direct evidence for synchrotron radiation, caused by relativistic electrons spiralling along magnetic fields, which naturally produces linearly polarized light. Measuring both the degree and orientation of polarization reveals the strength and structure of magnetic fields in jets and lobes, as well as information about the galaxy’s orientation relative to the observer.To quantitatively describe the polarization state of electromagnetic radiation, a set of four values is utilized known as Stokes parameters: I, Q, U, and V. These parameters are experimentally convenient as they correspond to sums or differences of measurable intensities, allowing for direct determination from observations. The complex linear polarization is defined as:
@@
The study of the polarization states of radio emissions enables us to analyse two important effects: Faraday rotation and Faraday depolarization (Anderson et al., 2015; Piras, 2024). Faraday rotation is caused by the rotation of the plane of polarization of electromagnetic waves as they journey across vast cosmic distances interacting with magnetized medium. Crucially, the extent of this rotation directly encodes vital information about the magneto-ionic medium—the ionized gas and embedded magnetic fields. This makes Faraday rotation an exceptionally powerful tool for probing magnetic fields not just within galaxies, but also in the vast and tenuous intergalactic medium, providing unique insights into the large-scale magnetization of the Universe that cannot be obtained through other means (Brentjens and de Bruyn, 2005). To accurately characterize the Faraday rotation for these distinct regions of polarized emission, the concept of Faraday depth is utilized. Faraday depth is defined as the integral of thermal electron density (cm-3) and parallel magnetic field (µG) from the observer to a point along the line of sight, with positive towards the source (Ferrière et al., 2021):
Within an extended source, emission at different physical depths contributes polarized radiation with distinct , enabling multi-component Faraday spectra even without beam depolarization.
For a simple foreground screen (all emission at screen), the observed rotation measure equals the screen Faraday depth. In general, is the first moment of the Faraday dispersion function , while traces the full depth structure. We use both quantities interchangeably when referring to a dominant single-screen case. In such an ideal case, the observed polarization angle is directly determined by the source’s intrinsic polarization angle and the Faraday rotation measure (RM), following the simple linear relationship (Anderson et al., 2016):
@@ 3 Depolarization
Depolarization occurs because variations in the magnetic field or electron density along the line of sight, or across the telescope beam, cause the polarization angle to rotate by different amounts in different regions, reducing the net observed polarization (Burn, 1966; Goodlet et al., 2004). When these diverse polarization orientations are averaged across the observing beam, they tend to cancel each other out, diminishing the overall polarization signal. Hence, two types of depolarization come into play: internal and external (Pasetto et al., 2018). The remaining polarization percentage is effectively utilized to gauge the degree of order and orientation within the magnetic field, whether it resides in the radio source itself or in the intervening inter galactic medium (IGM). The complex polarized signal observed from radio sources is fundamentally described by a model that accounts only for Faraday rotation and does not depolarize the signal is given as (Burn, 1966) :
Here, represents the intrinsic fractional polarization of the radiation, and is the intrinsic polarization angle at the source of the emission. The term describes the Faraday rotation caused by the foreground magneto-ionic medium, where RM is the rotation measure and is the square of the observing wavelength. This model is also referred as ‘Faraday thin’ (Burn, 1966). In real scenarios, the magnetic field is not perfectly ordered, there exists random and uniform fields causing depolarization (Goodlet et al., 2004). The various depolarization contributions are thus added to this fundamental equation 5.
@@
Note, this implementation of the Faraday-thick (Burn-slab) model omits a foreground Faraday rotation term (RM_fg), following Sokoloff et al. (1998) and as implemented in the version of RM-Tools used here. For instrumental leakage near RM 0, this approximation is reasonable, as ionospheric RM shifts are small ( rad m-2). However, for astrophysical components, this may bias the recovered Faraday depth if a significant foreground screen is present; we discuss this further in Section 6.
Moreover, for an external Faraday dispersion component, the fractional polarization scales as,
(Burn, 1966), so that even intrinsically 100% polarized emission is exponentially suppressed once . For the LoTSS HBA band (120–168 MHz, ), this implies that components with are already heavily depolarized across most of the band, and for they would be effectively undetectable at LoTSS sensitivity even if intrinsically strongly polarized.
@@ 4 The LoTSS DR2 Catalogue and the Selected Source
The present work is explicitly designed as a pilot for a larger qu-fitting campaign on the LoTSS-DR2 FR-I like population. For this purpose, a single source is preferred to satisfy three criteria: (i) it should be representative of the catalogue in terms of luminosity and size, (ii) it must have a robust polarization and RM measurement in DR2, and (iii) its angular and projected linear size should be large enough that depolarization effects due to propagation through an extended magneto-ionic medium are expected to be significant.
In this pilot study, we therefore focus on ILTJ012215.21-254334.8, which satisfies these criteria. It has an estimated luminosity of W Hz-1 at 144 MHz, RM of rad m-2, and a projected linear size of 335 kpc, i.e., very close to the LoTSS-DR2 median linear size of 300 kpc (O’Sullivan et al., 2023). he source has an angular size of , comparable to the 20” LoTSS beam, so the extracted spectrum averages over the entire source. With an estimated luminosity of at 144 MHz (below the classical FR break of ), this is classified as an FR-I radio galaxy (Fanaroff and Riley, 1974). The polarization fraction of approximately and RM are typical of the catalogued sources, so it can be regarded as representative of the dominant population of low-luminosity FR-I-like radio galaxies in LoTSS-DR2. By analysing this object in detail, the intention is to (a) validate the qu-fitting setup, priors, and instrumental-leakage treatment under realistic LoTSS conditions, and (b) identify which depolarization parameters and model classes are actually constrained by the data for a typical source. These lessons will directly inform the design and interpretation of a forthcoming study of the full DR2 FR-I sample.
@@
| Morphology | FR-I |
Here, RM is the Rotation Measure; P is the linear polarization; is the fractional polarization; I is the total intensity; is the luminosity at 144 MHz; is the projected linear size; and z is the redshift. The 335 kpc projected size implies a substantial path length through the host and cluster magneto-ionic medium, enhancing sensitivity to line-of-sight Faraday gradients and making multi-component depolarization more likely than for compact sources.
[…]
@@ 5.2 The priors
In qu-fitting, setting good starting assumptions (priors) for the model parameters is important because they help guide the fitting process in the right direction. The selection of these priors is often based on the observed Faraday spectrum of the source (Piras, 2024). For example, when combining multiple models, The total initial polarization fraction across all astrophysical components must satisfy . The instrumental leakage term is excluded from this physical constraint, and each initial polarization angle is limited between 0 and 180 degrees for all the models considered in Table 2.
@@
To analyse the observed Stokes and values, along with their uncertainty, a one-dimensional qu-fitting technique was applied to the LoTSS DR2 polarization spectra of the source111https://lofar-mksp.org/data/. The LoTSS DR2 RM Grid lacks reliable channelized Stokes I spectra222https://github.com/sposullivan/LoTSS-RM-Grid/blob/main/example_RMtable_PolSpectra.ipynb, so we construct a model from the catalogue 144 MHz flux density assuming a spectral index . The Stokes I uncertainty is taken as the catalogue measurement uncertainty, held constant across all frequency channels.
The fractional uncertainties are then computed as:
This conservative approach ensures that and Bayes factor model selection are driven by the spectral shape of and , making the analysis robust to absolute polarization scale uncertainties while remaining sensitive to depolarization signatures.
We used the RM-TOOLS software333https://github.com/CIRADA-Tools/RM-Tools/tree/master (Purcell et al., 2020), which incorporates the MultiNest algorithm (Feroz and Hobson, 2008; Feroz et al., 2009, 2019). MultiNest, implemented via PYMULTINEST (Buchner, J. et al., 2014), is employed to efficiently explore the parameter space of the chosen model, identify the best-fit parameters, and estimate their associated uncertainties. The algorithm provides the posterior distribution for the model parameters and the Bayesian evidence () for the model, which quantifies the probability of observing the data given a specific model.
@@
Model comparison follows the Kass and Raftery (1995) scale for , where values are barely worth mentioning, values between and indicate positive evidence, values between and indicate strong evidence, and values indicate very strong evidence against the alternative. Thus, decisively favours the reference model. Here, indicates very strong evidence for over alternatives (Kass and Raftery, 1995)).
@@ 6 Results
The comparison of different models based on the odds ratio and the reduced chi-squared statistic () indicates that the best description of the depolarization behaviour of the source is provided by the 1T+2ED model. This particular model includes three polarization components: one Faraday-thin term corresponding to instrumental leakage, located at a rotation measure (RM) of about , and two external depolarization (2ED) components, centred near , which likely represent the astrophysical polarized emission (O’Sullivan et al., 2023). The astrophysical emission is better represented by the superposition of two external depolarization components at similar RMs, each capturing slightly different aspects of turbulence or inhomogeneities in the magneto-ionic medium along the line of sight (Sokoloff et al., 1998; O’Sullivan et al., 2017).
@@
| 1T+2ED | 11 | 3 | 814 | 2.12 | 1385.46 | 0 |
| 1Tk+2ED | 11 | 3 | 814 | 2.12 | 1377.38 | 16.16 |
| 1T+1ED | 7 | 2 | 818 | 2.77 | 1154.35 | 462.23 |
| 1Tk+1ED | 7 | 2 | 818 | 2.77 | 1148.94 | 473.05 |
| 1ED | 4 | 1 | 821 | 2.94 | 1097.37 | 576.18 |
Note here the Faraday-thick (Burn-slab) model omits a foreground RMfg term (Sokoloff et al., 1998), where effective RM is tied to dispersion origin. This limits flexibility for ionospheric-shifted instrumental leakage. The modest degeneracy 1T+2ED () vs 1Tk+2ED () partly reflects this limitation.
To test this concern directly, we re-implemented RMfg in thick models (1Tk+1ED, 1Tk+2ED) with RM (free effective RM) and prior RM (O’Sullivan et al., 2023).
Results confirm 1T+2ED remains decisively preferred even with equal model flexibility: 1Tk+2ED+RMfg (; positive evidence against), 1T+1ED (; very strong evidence against), and 1Tk+1ED+RMfg (; very strong evidence against).
Thin leakage + 2ED is physically preferred (generally Bayes factors favour simpler model following Occam’s Razor). Astrophysical components (RM –) remain unaffected by ionospheric modelling. We therefore retain the standard Sokoloff form (Sokoloff et al., 1998) for thick models, ensuring consistency with previous low-frequency studies.
@@
The narrow astrophysical component (, RM rad m-2, rad m-2) has negligible dispersion (consistent with ), resembling a foreground Faraday-thin screen. It could equivalently be modelled as thin, but allowing dispersion improves the fit near m2 without over-fitting (Bayes factor remains decisive).
The broad component (, RM rad m-2, rad m-2) is separated in RM-space by rad m, suggesting distinct magneto-ionic volumes (e.g., foreground cloud + lobe plasma). Its dispersion depolarizes it by at m2 but at m2, explaining why it dominates only high-frequency channels (hence absent in 1T+1ED fits).
The narrow component carries of total but dominates mid-band signal, placing an upper limit on undetected thicker emission ( rad m-2, invisible in LoTSS due to strong depolarization effects).
Beyond confirming that more than one external Faraday dispersion component is preferred, the fits yield several quantitative constraints that are directly relevant for future population studies (Table 4). First, the two astrophysical components are tightly clustered in Faraday depth around RM rad m-2 while requiring distinct dispersion widths and intrinsic polarization fractions, indicating at least two magneto-ionic regions along the line of sight that are co-located in RM but differ in turbulence level or covering factor. Second, the relative amplitudes of these components imply that the more strongly dispersed screen contributes the majority of the depolarization across the LoTSS band, whereas the narrower component remains detectable even at the longest wavelengths, providing a concrete example of how LoTSS preferentially selects less-depolarized sub-components within otherwise complex sources.
@@
To investigate the elevated , we examined , residuals normalized by their measurement uncertainties, as a function of . Figure 3 shows noise-like residuals without strong frequency correlations, favouring underestimated noise over un-modelled Faraday structure. The discrepancy likely combines moderate model limitations with noise underestimation typical of interferometric imaging.
Figure 4 presents the de-rotated polarization angle , where the LoTSS-DR2 catalogue RMcat = -47.15 rad m-2 has been used for de-rotation. A single thin Faraday screen at this RM would produce a flat line. Instead, the data exhibit significant scatter and curvature around this reference.
Linear fitting yields a residual rotation measure RMres = +0.28 6.99 rad m-2, statistically consistent with the catalogue value (consistent with dominant foreground screen removal). However, the quadratic term is highly significant (), indicating substantial departure from simple Faraday-thin screen behaviour.
The reduced further confirms systematic deviations from linear behaviour, characteristic of either multiple discrete Faraday-thin screens along the line-of-sight, or Faraday dispersion within an extended magneto-ionic medium. This complexity also motivates our adoption of the 1T+2ED Faraday dispersion function modelling presented in Section 3.
Figure 5 illustrates depolarization via vector contraction in the - plane. High-frequency (blue) points cluster near the outer radius, indicating strong, coherent polarization. Low-frequency (red) points cluster near the centre, indicating severe depolarization. This trend is the same depolarization seen in Figure 2, but represented geometrically rather than via polarized intensity vs .
Finally, the posterior distribution of the best-fit model parameters along with error bars is shown in Figure 6. Note that the derived values of in Table 4 are not meaningful due to the lack of absolute calibration for the polarization angles.